OpenStudy (anonymous):

ms.Lewis plans to drive 900 miles to her vacation destination, driving an average of 50 miles per hour. How many miles per hour faster must she average, while driving to reduce her total driving time by 3 hours? A. 5 B. 8 C. 10 D. 15 E. 18

OpenStudy (zpupster):

distance = rate*time 900=50*t 900/50=t this is the time it takes her at 50mph, so take this t and subtract 3 and place back in formula 900=r*t 900/t=r r= the rate it takes at 3 hours less lastly take the r here r-50=rate it will take to cut driving time by 3 hours